# Pimba: A Processing-in-Memory Acceleration for Post-Transformer Large Language Model Serving

### v1

@@ Abstract.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 58th IEEE/ACM International Symposium on Microarchitecture; October 18–22, 2025; Seoul, Koreaisbn: 978-X-XXXX-XXXX-X/XX/XX
@@ 1. Introduction
To evaluate Pimba’s effectiveness, we use four post-transformer LLM models–Mamba-2, GLA, RetNet, and HGRN2 (Dao and Gu, 2024; Yang et al., 2024b; Sun et al., 2023; Qin et al., 2024)– along with Zamba2, a hybrid transformer-Mamba-2 model (Zyphra, [n. d.]), and OPT (Zhang et al., 2022), a traditional attention-based model. Our experimental results show that compared to LLM-optimized GPU and GPU+PIM systems, Pimba achieves 14.6$\times$ and 6.9$\times$ lower latency in state update operations, resulting in up to 4.1$\times$ and 2.1$\times$ higher throughput, respectively, with minimal area overhead on the memory device. These advantages in both performance and area-efficiency demonstrate that Pimba is an effective PIM-based solution for LLM serving, capable of supporting both transformer and post-transformer models, paving the way toward scalable and cost-efficient deployment of emerging LLM architectures.
@@ 3.2. Quantization Analysis for SU-LLMs
Transformer-based LLMs exhibit negligible perplexity increases across all formats, while SU-LLMs exhibit a severe increase in perplexity with floating point formats (e.g. 8,126 for GLA with e4m3). This discrepancy arises from the SU-LLMs’ continuous state “update” mechanism. This makes them vulnerable to loss of small values during accumulation due to limited mantissa precision, which is called swamping effect (Higham, 1993; Wang et al., 2018). The 7-bit (int8) and 6-bit (MX8) mantissas provide enough precision to mitigate swamping, whereas the 3-bit and 2-bit mantissas in e4m3 and e5m2 render these formats highly susceptible. This finding aligns with conventional practices in training deep learning models, wherein weights are stored at higher precisions to reduce numerical errors (Micikevicius et al., 2018). Another notable observation is that stochastic rounding has a substantial positive impact on SU-LLMs, in contrast to transformer-based LLMs. For example, the perplexity of Mamba-2 in the e5m2 format drops dramatically from 62 to 12.2 when stochastic rounding is applied. In SU-LLMs, stochastic rounding probabilistically preserves smaller magnitude values that would otherwise be lost due to swamping, thereby maintaining more information during state update scenario.
@@ 5.1. Overview
1
Pimba system. Pimba handles user requests in two phases: prefill and generation. In the prefill phase, all operations, including state updates, run on the GPU, as they can be restructured into compute-intensive forms (Sun et al., 2023; Dao and Gu, 2024). In the generation phase, Pimba offloads the state update and attention operations to the PIM, while other operations remain on the GPU. For the PIM-executed operations, Pimba transfers operands to PIM registers, computes partial sums, and sends the results back to the GPU for accumulation.
@@
2
PIM architecture. Pimba aims to accelerate state updates and attention while minimizing PIM area overhead. To achieve its objectives, Pimba introduces two key architectural innovations: (1) Pimba employs a novel State-update Processing Unit (SPU) that is shared between two memory banks (Section 5.2). Unlike traditional designs where a processing unit can only access one bank at a time for either read or write (Lee et al., 2021), Pimba’s pipelined design allows simultaneous reading from one bank and writing to the other. This overlapping of read and write operations across two banks enables Pimba to halve the number of processing units compared to a per-bank design, while maintaining the same throughput. (2) Within each SPU, Pimba integrates an MX-based State-update Processing Engine (SPE) to enable area-efficient and accuracy-preserving computations (Section 5.3). While prior works primarily focus on MX quantization for dot product operations (Kim et al., 2024), Pimba introduces a microarchitectural design specifically tailored for MX-based element-wise addition and multiplication.
@@
3
PIM banks with state data layout. We divide each state column along the $dim_{head}$ dimension into sub-chunks based on the DRAM column size. Then, we group sub-chunks across the $dim_{state}$ dimension to form a chunk that aligns with the DRAM row size, enabling efficient sequential access within each SPU. To further maximize operand reuse across chunks, we organize the chunks into chunk groups and assign each group to a DRAM bank. Chunks within the same group share the operands $d_{t}$, $q_{t}$, and $k_{t}$, and are placed in consecutive rows of the bank. This arrangement enables the transfer of shared operands to Pimba once per chunk group, while only the corresponding $v_{t}$ vector is transferred per chunk, enhancing data reuse. Once assigned to banks, Pimba processes sub-chunks sequentially in a pipelined manner by reading consecutive columns within a DRAM row. In the $i$-th iteration (time unit during which Pimba processes each sub-chunk), Pimba utilizes the $i$-th sub-chunk, the shared $d_{t}$, $q_{t}$, $k_{t}$ vectors, and the $i$-th element of the $v_{t}$ vector in its computations.
@@ 5.4. Attention Support in Pimba
The attention computation in Pimba proceeds in two phases: score and attend. In the score phase, Pimba performs dot products between the query vector and key vectors using its in-pipeline dot product unit as shown in the score dataflow of Figure 10(b). The intermediate results are sent to the GPU, where they are accumulated and passed through a softmax function to produce normalized attention scores. In the attend phase, Pimba multiplies each attention score with its corresponding value vector and accumulates the results. This phase utilizes Pimba’s multiplier and adder units, as shown in the attend dataflow of Figure 10. By leveraging a shared microarchitectural substrate, Pimba achieves efficient support for both state update and attention operations, demonstrating the versatility of its PIM-based execution model.
@@ 6.1. Methodology
RetNet GPU 15.84 72.2 44.0 42.0 59.6 26.0 53.4 47.2
Pimba 15.95 72.2 43.7 42.0 59.7 25.7 53.2 47.0 (–0.2)
GLA GPU 15.53 71.7 43.6 41.8 59.2 26.5 54.9 47.3
Pimba 15.58 71.8 43.5 41.9 59.0 26.6 56.0 47.5 (+0.2)
HGRN2 GPU 14.48 73.0 48.5 44.6 60.6 24.9 54.6 48.5
Pimba 15.09 73.0 48.7 44.9 60.8 25.6 54.7 48.8 (+0.3)
Mamba-2 GPU 11.46 76.2 59.7 49.6 69.7 33.1 64.2 56.7
Pimba 11.51 76.1 59.1 49.6 69.5 33.7 63.8 56.7 (+0.0)
Zamba2 GPU 5.94 78.9 65.0 63.8 78.6 53.9 77.3 68.9
Pimba 5.96 78.5 64.7 63.8 78.7 53.8 77.0 68.7 (–0.2)
OPT GPU 12.29 76.3 63.3 50.5 65.6 30.5 65.3 56.3
Pimba 12.29 76.1 63.1 50.5 65.7 30.7 65.0 56.3 (+0.0)
@@
6.2. Results
Throughput. Figure 12 reports the normalized token-generation throughput across various SU-LLMs and an attention-based LLM using (2,048, 2,048) input/output sequence lengths. The GPU+Q system achieves an average of 1.4$\times$ higher throughput over the GPU baseline due to halving the state size. Interestingly, GPU+PIM also achieves 1.4$\times$ throughput improvement over GPU, matching or occasionally underperforming compared to GPU+Q despite leveraging PIM. This occurs because HBM-PIM employs a time-multiplexed design without Pimba’s access interleaving, causing state update operations to span multiple cycles. Consequently, even with PIM, internal bandwidth is underutilized, yielding sub-optimal performance. Furthermore, GPU+PIM processes twice as much data compared to GPU+Q, as it uses fp16 states rather than 8-bit formats. In contrast, Pimba consistently outperforms GPU and GPU+PIM, providing average throughput gains of 2.0$\times$ and 1.4$\times$ (up to 4.1$\times$ and 2.1$\times$), respectively. This superior performance results from efficiently leveraging internal bandwidth through pipelining and access interleaving, alongside the advantages of low-precision state representation.
Latency breakdown. To identify the main contributors to throughput improvements, Figure 13 presents normalized latency for large scale SU-LLMs and an attention-based LLM during the generation phase. As shown, Pimba reduces state update operation latency by 14.6$\times$ and 6.9$\times$ compared to GPU and GPU+PIM, respectively, highlighting its efficient handling of state updates. Additionally, the results show greater overall latency reduction for larger batch sizes and models dominated by state update operations. For instance, HGRN2 with a batch size of 32 spends only 14% of its latency on state updates, resulting in a modest 1.2$\times$ latency reduction. In contrast, RetNet with a batch size of 128 allocates 74% of its latency to state updates, achieving a substantial 3.2$\times$ latency reduction. Moreover, Pimba achieves latency reductions for attention operations by 6.3$\times$ and 1.8$\times$ compared to GPU and GPU+PIM, respectively, effectively accelerating attention-dominated hybrid models such as Zamba2 and the attention-based OPT. Notably, the latency reduction for attention operations in Pimba is smaller compared to state updates. This is due to attention operations primarily consisting of GEMV computations without requiring write operations, thus limiting the benefits of Pimba’s pipelined design and access interleaving. Nevertheless, by employing the MX8 format, Pimba achieves a 1.8$\times$ attention latency reduction over GPU+PIM, demonstrating robust performance in hybrid models as well.
Energy consumption. Figure 14 presents the normalized energy consumption breakdown of large scale SU-LLMs and an attention-based LLM with a batch size of 128 during the generation phase. The results show that Pimba achieves an average of 2.2$\times$ lower energy consumption compared to GPU, demonstrating its cost-effectiveness for SU-LLM serving. These improvements stem from reducing the amount of state and KV cache transfer between the GPU and memory while performing state update operations within the PIM itself. Additionally, Pimba exhibits an average of 1.2$\times$ lower energy consumption compared to GPU+PIM, attributed to Pimba’s use of the MX8 which significantly reduces transfer between the GPU as well as computation time.
@@ 6.2. Results
Area. To assess the practicality of Pimba, we measure the area and power consumption of Pimba and HBM-PIM. Table 3 presents the results. For a fair comparison, HBM-PIM is optimized for state update computations by retaining only the essential components while reducing or removing others, such as shrinking internal registers and eliminating specific control logic. As indicated in the table, Pimba adheres to PIM area constraints, with an area overhead of 13.4%, well below the 25% maximum logic ratio recommended by prior work (He et al., 2020). Although Pimba incurs an area overhead approximately 1.6% larger than HBM-PIM, this increase is justified by delivering up to a 2.1$\times$ throughput improvement over HBM-PIM.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: 58th IEEE/ACM International Symposium on Microarchitecture; October 18–22, 2025; Seoul, Republic of Koreabooktitle: 58th IEEE/ACM International Symposium on Microarchitecture (MICRO ’25), October 18–22, 2025, Seoul, Republic of Koreadoi: 10.1145/3725843.3756121isbn: 979-8-4007-1573-0/2025/10ccs: Computer systems organization Neural networksccs: Computer systems organization Heterogeneous (hybrid) systems
@@ 1. Introduction
To evaluate Pimba’s effectiveness, we use four post-transformer LLM models–Mamba-2, GLA, RetNet, and HGRN2 (Dao and Gu, 2024; Yang et al., 2024b; Sun et al., 2023; Qin et al., 2024)– along with Zamba2, a hybrid transformer-Mamba-2 model (Zyphra, [n. d.]), and OPT (Zhang et al., 2022), a traditional attention-based model. Our experimental results show that compared to LLM-optimized GPU and GPU+PIM systems, Pimba achieves 14.6$\times$ and 6.9$\times$ lower latency in state update operations, resulting in up to 4.1$\times$ and 2.1$\times$ higher throughput, respectively, with minimal area overhead on the memory device. These advantages in both performance and area-efficiency demonstrate that Pimba is an effective PIM-based solution for LLM serving, capable of supporting both transformer and post-transformer models, paving the way toward scalable and cost-efficient deployment of emerging LLM architectures. A full-system simulator for Pimba and the accuracy evaluation code are open-sourced at https://github.com/casys-kaist/pimba.
@@ 3.2. Quantization Analysis for SU-LLMs
Transformer-based LLMs exhibit negligible perplexity increases across all formats, while SU-LLMs exhibit a severe increase in perplexity with floating point formats (e.g. 8,114 for GLA with e4m3). This discrepancy arises from the SU-LLMs’ continuous state “update” mechanism. This makes them vulnerable to loss of small values during accumulation due to limited mantissa precision, which is called swamping effect (Higham, 1993; Wang et al., 2018). The 7-bit (int8) and 6-bit (MX8) mantissas provide enough precision to mitigate swamping, whereas the 3-bit and 2-bit mantissas in e4m3 and e5m2 render these formats highly susceptible. This finding aligns with conventional practices in training deep learning models, wherein weights are stored at higher precisions to reduce numerical errors (Micikevicius et al., 2018). Another notable observation is that stochastic rounding has a substantial positive impact on SU-LLMs, in contrast to transformer-based LLMs. For example, the perplexity of Mamba-2 in the e5m2 format drops dramatically from 62 to 11.9 when stochastic rounding is applied. In SU-LLMs, stochastic rounding probabilistically preserves smaller magnitude values that would otherwise be lost due to swamping, thereby maintaining more information during state update scenario.
@@ 5.1. Overview
(1) Pimba system. Pimba handles user requests in two phases: prefill and generation. In the prefill phase, all operations, including state updates, run on the GPU, as they can be restructured into compute-intensive forms (Sun et al., 2023; Dao and Gu, 2024). In the generation phase, Pimba offloads the state update and attention operations to the PIM, while other operations remain on the GPU. For the PIM-executed operations, Pimba transfers operands to PIM registers, computes partial sums, and sends the results back to the GPU for accumulation.
@@
(2) PIM architecture. Pimba aims to accelerate state updates and attention while minimizing PIM area overhead. To achieve its objectives, Pimba introduces two key architectural innovations: (1) Pimba employs a novel State-update Processing Unit (SPU) that is shared between two memory banks (Section 5.2). Unlike traditional designs where a processing unit can only access one bank at a time for either read or write (Lee et al., 2021), Pimba’s pipelined design allows simultaneous reading from one bank and writing to the other. This overlapping of read and write operations across two banks enables Pimba to halve the number of processing units compared to a per-bank design, while maintaining the same throughput. (2) Within each SPU, Pimba integrates an MX-based State-update Processing Engine (SPE) to enable area-efficient and accuracy-preserving computations (Section 5.3). While prior works primarily focus on MX quantization for dot product operations (Kim et al., 2024), Pimba introduces a microarchitectural design specifically tailored for MX-based element-wise addition and multiplication.
@@
(3) PIM banks with state data layout. We divide each state column along the $dim_{head}$ dimension into sub-chunks based on the DRAM column size. Then, we group sub-chunks across the $dim_{state}$ dimension to form a chunk that aligns with the DRAM row size, enabling efficient sequential access within each SPU. To further maximize operand reuse across chunks, we organize the chunks into chunk groups and assign each group to a DRAM bank. Chunks within the same group share the operands $d_{t}$, $q_{t}$, and $k_{t}$, and are placed in consecutive rows of the bank. This arrangement enables the transfer of shared operands to Pimba once per chunk group, while only the corresponding $v_{t}$ vector is transferred per chunk, enhancing data reuse. Once assigned to banks, Pimba processes sub-chunks sequentially in a pipelined manner by reading consecutive columns within a DRAM row. In the $i$-th iteration (time unit during which Pimba processes each sub-chunk), Pimba utilizes the $i$-th sub-chunk, the shared $d_{t}$, $q_{t}$, $k_{t}$ vectors, and the $i$-th element of the $v_{t}$ vector in its computations.
@@ 5.4. Attention Support in Pimba
The attention computation in Pimba proceeds in two phases: score and attend. In the score phase, Pimba performs dot products between the query vector and key vectors using its in-pipeline dot product unit as shown in the score dataflow of Figure 10(b). The intermediate results are sent to the GPU, where they are accumulated and passed through a softmax function to produce normalized attention scores. In the attend phase, Pimba multiplies each attention score with its corresponding value vector and accumulates the results. This phase utilizes Pimba’s multiplier and adder units, as shown in the attend dataflow of Figure 10(b). By leveraging a shared microarchitectural substrate, Pimba achieves efficient support for both state update and attention operations, demonstrating the versatility of its PIM-based execution model.
@@ 6.1. Methodology
RetNet GPU 15.83 72.3 44.0 42.0 59.5 25.5 53.1 47.0
Pimba 15.95 72.3 43.7 41.9 59.7 25.8 53.7 47.1 (+0.1)
GLA GPU 15.54 71.6 43.8 41.8 59.1 26.7 55.4 47.5
Pimba 15.57 71.5 43.4 41.9 59.1 26.7 55.2 47.4 (–0.1)
HGRN2 GPU 14.48 73.1 48.5 44.6 60.7 25.3 54.7 48.6
Pimba 15.09 73.0 48.5 44.8 60.3 25.2 55.0 48.6 (+0.0)
Mamba-2 GPU 11.46 76.4 59.6 49.6 69.4 33.2 64.0 56.7
Pimba 11.51 76.3 59.4 49.6 69.3 33.2 63.9 56.6 (–0.1)
Zamba2 GPU 5.94 78.9 64.9 63.8 78.9 53.8 77.7 69.0
Pimba 5.96 78.6 64.4 63.7 78.4 53.8 77.1 68.7 (–0.3)
OPT GPU 12.29 76.2 63.3 50.5 65.6 30.6 65.1 56.3
Pimba 12.29 76.1 63.2 50.5 65.6 30.3 65.2 56.2 (–0.1)
6.2. Results
Throughput. Figure 12 reports the normalized token-generation throughput across various SU-LLMs and an attention-based LLM using (2,048, 2,048) input/output sequence lengths. The GPU+Q system achieves an average of 1.4$\times$ higher throughput over the GPU baseline due to halving the state size. Interestingly, GPU+PIM also achieves 1.4$\times$ throughput improvement over GPU, matching or occasionally underperforming compared to GPU+Q despite leveraging PIM. This occurs because HBM-PIM employs a time-multiplexed design without Pimba’s access interleaving, causing state update operations to span multiple cycles. Consequently, even with PIM, internal bandwidth is underutilized, yielding sub-optimal performance. Furthermore, GPU+PIM processes twice as much data compared to GPU+Q, as it uses fp16 states rather than 8-bit formats. In contrast, Pimba consistently outperforms GPU and GPU+PIM, providing average throughput gains of 1.9$\times$ and 1.4$\times$ (up to 4.1$\times$ and 2.1$\times$), respectively. This superior performance results from efficiently leveraging internal bandwidth through pipelining and access interleaving, alongside the advantages of low-precision state representation.
Latency breakdown. To identify the main contributors to throughput improvements, Figure 13 presents normalized latency for large scale SU-LLMs and an attention-based LLM during the generation phase. As shown, Pimba reduces state update operation latency by 14.6$\times$ and 6.9$\times$ compared to GPU and GPU+PIM, respectively, highlighting its efficient handling of state updates. Additionally, the results show greater overall latency reduction for larger batch sizes and models dominated by state update operations. For instance, HGRN2 with a batch size of 32 spends only 14% of its latency on state updates, resulting in a modest 1.2$\times$ latency reduction. In contrast, RetNet with a batch size of 128 allocates 74% of its latency to state updates, achieving a substantial 3.2$\times$ latency reduction. Moreover, Pimba achieves latency reductions for attention operations by 6.3$\times$ and 2.1$\times$ compared to GPU and GPU+PIM, respectively, effectively accelerating attention-dominated hybrid models such as Zamba2 and the attention-based OPT. Notably, the latency reduction for attention operations in Pimba is smaller compared to state updates. This is due to attention operations primarily consisting of GEMV computations without requiring write operations, thus limiting the benefits of Pimba’s pipelined design and access interleaving. Nevertheless, by employing the MX8 format, Pimba achieves a 2.1$\times$ attention latency reduction over GPU+PIM, demonstrating robust performance in hybrid models as well.
@@ 6.2. Results
Energy consumption. Figure 14 presents the normalized energy consumption breakdown of large scale SU-LLMs and an attention-based LLM with a batch size of 128 during the generation phase. The results show that Pimba achieves an average of 2.2$\times$ lower energy consumption compared to GPU, demonstrating its cost-effectiveness for SU-LLM serving. These improvements stem from reducing the amount of state and KV cache transfer between the GPU and memory while performing state update operations within the PIM itself. Additionally, Pimba exhibits an average of 1.3$\times$ lower energy consumption compared to GPU+PIM, attributed to Pimba’s use of the MX8 which significantly reduces transfer between the GPU as well as computation time.
@@
Area. To assess the practicality of Pimba, we measure the area and power consumption of Pimba and HBM-PIM. Table 3 presents the results. For a fair comparison, HBM-PIM is optimized for state update computations by retaining only the essential components while reducing or removing others, such as shrinking internal registers and eliminating specific control logic. As indicated in the table, Pimba adheres to PIM area constraints, with an area overhead of 13.4%, well below the 25% maximum logic ratio recommended by prior work (He et al., 2020). Although Pimba incurs an area overhead approximately 1.5% larger than HBM-PIM, this increase is justified by delivering up to a 2.1$\times$ throughput improvement over HBM-PIM.
@@ 9. Conclusion
Acknowledgements.
We thank the anonymous reviewers for their comments and feedback. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) (No.RS-2024-00396013), IITP under the Graduate School of Artificial Intelligence Semiconductor (IITP-2025-RS-2023-00256472), grant funded by the Korea government (MSIT). This work was also supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by ICT R&D program of MSIT/IITP (No.RS-2025-02305453, Development of distributed inference and model optimization technology for heterogeneous Al semiconductors). The EDA tool was supported by the IC Design Education Center (IDEC), Korea.